MATEMATICA Anno Accademico 2005-2006 Scienze Geologiche M.Novaga, V.M. Tortorelli

————————————————————————————————————————–

II prova in itinere, 30 gennaio 2006 RISULTATI

————————————————————————————————————————–

ALEARDI Mattia  discreto
BACHINI Viola  suffi
BELLONI Rossella  vale
BIONDI Giacomo  vale
GIANELLA Valentina  suffi
IACCARINO Salvatore  vale+
MARCHETTI Samuele  vale
NIRTA Antonio  ins
PAOLINI Marco  vale
PARDINI Dario  vale+
PELLEGRINI Andrea  ins
PIERETTI Carlo  ins
SECCHIARI Arianna  ins
TAMBELLINI Simone  vale+
VITI Federico  vale